# AutoPrep: Natural Language Question-Aware Data Preparation with a Multi-Agent Framework

### v3

@@ Abstract.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
@@
The source code, data, and/or other artifacts have been made available at https://github.com/fmh1art/AutoPrep.
@@ 1. Introduction
Tabular Question Answering (TQA) refers to the task of answering natural language (NL) questions based on provided tables (Jin et al., 2022; Cheng et al., 2022, 2023; Pal et al., 2023). TQA empowers non-technical users such as domain scientists to easily analyze tabular data and has a wide range of applications, including table-based fact verification (Chen et al., 2019; Aly et al., 2021) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022). As TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024; Wang et al., 2024) rely on large language models (LLMs).
@@
(3) Irrelevant Columns. In some tables, although many columns are available, only a few are actually relevant to the NL question. For example, as shown in Figure 2(c), $Q_{7}$ does not require the Summary column in table $T_{7}$. Similarly, in $T_{8}$, although there are 12 columns, only 2 are needed to answer $Q_{8}$. Thus, column filtering, i.e., retaining only the columns directly related to the question, is crucial for TQA, as irrelevant columns add complexity and can mislead LLMs into producing incorrect answers. For instance, the Summary column in $T_{7}$ contains an average of 100 words, increasing the difficulty for LLMs to identify the relevant AirDate column.
This highlights the critical role of thorough data prep in ensuring accurate TQA. Note that, although issues like missing values and duplicates are common in real-world data, they are less prevalent than the three highlighted issues in current TQA benchmarks. Nevertheless, we recognize their significance and leave the exploration of broader data prep challenges to future work.
@@
AutoPrep: A Hierarchical, Multi-Stage Approach. To address the challenges, we propose AutoPrep, which features two key ideas. First, drawing inspiration from modern DBMS, particularly the distinction between logical operations and their physical implementations, AutoPrep separates high-level, logical data prep operations from the concrete methods used for execution. Specifically, it introduces a planning stage that generates a logical plan for each question, consisting of a sequence of high-level data prep operations tailored to the question’s needs, such as column derivation, normalization and filtering shown in Figure 2. In the next stage, AutoPrep maps these logical operations to the corresponding physical implementations. This separation allows AutoPrep to break down complex data prep tasks into smaller, more manageable sub-tasks, making the process more easier to solve. Moreover, this modular design makes AutoPrep extensible, enabling the development of specialized implementations for each type of data prep operation or the introduction of new operation types.
@@ 2.1. Tabular Question Answering
Figure 2 provides several examples of TQA. For instance, consider table $T_{1}$, which contains medal information for cyclists from different countries, with two columns: Cyclist and Medal. Given the question $Q_{1}$, “Which country has the most medals?”, the answer should be ITA, as two Italian cyclists, “Dav” and “Alex”, have won a total of 4 medals, more than the ESP cyclist “Alej”.
TQA often requires complex reasoning over tables. For instance, to answer question $Q_{2}$, we first need to calculate the “GDP growth rate” for all countries, then sort the countries by growth rate, and finally identify the country with the highest GDP growth rate, i.e., CHN.
@@ 2.2. Data Prep for TQA
(a) The Derive operation: Figure 2(a) shows three examples of the Derive operations, i.e., extracting Country information from column Cyclist in $T_{1}$, computing new column GrowthRate using two columns in $T_{2}$, and inferring a status of IfLost by analyzing the scores in column Result. Note that the corresponding Python implementations are provided alongside the operations.
@@ 2.3. A Straightforward LLM-based Solution
First, at the logical-operation level, given the inherent difficulties in understanding both NL questions and tables, it is challenging to accurately identify which data prep operations are specifically required to satisfy the needs of the NL question. This often leads to false negatives (FNs) and false positives (FPs). For example, as shown in Figure 3, converting the Age column to a numerical format in code block ② is a false positive (FP), as it is irrelevant to the question. In contrast, failing to normalize the Date column before extracting the month in code block ③ constitutes a false negative (FN), as the method ignores the inconsistency in Date formats.
@@ 3. An Overview of AutoPrep
    Multiple Programmer Agents (i.e., Normalize): the Programming stage. It directs the LLM to generate physical implementation $f_{i}$ (e.g., Python code) for each operation $o_{i}$ customized for the table $T$. Besides, it is also tasked for code debugging if any execution errors occur.
@@
  • Normalize(“Format date as %m-%d”, Date) standardizes the values in the Date column into a unified format to support the ‘in February’ condition in the question.
@@
(b) The Programming stage: AutoPrep designs specialized Programmer agents for each operation type, i.e.,, Derive, Normalize and Filter. Each specialized Programmer focuses on generating executable code for its assigned operations.
@@
The Planner Agent. The key challenge is how to suggest a logical plan that address specific NL questions. Specifically, even for the same table, different NL questions may require not only different logical operations but also varying sequences of those operations. To address this challenge, we propose a novel Chain-of-Clauses (CoC) reasoning method for the Planner agent. This method translates the NL question into an Analysis Sketch, outlining how the table should be transformed to produce the answer, thereby guiding the agent’s reasoning based on this sketch. Compared to existing CoT methods (Wei et al., 2022), which break down questions into sub-questions, our approach better captures the relationships between questions and tables. More details of the method are given in Section 4.
@@ 4.1. A Direct Prompting Method
(b) Missing operations: Observing the ground-truth in Figure 4, we see that the Derive operation on Cyclist is not generated, as the column has already been filtered out. Moreover, although the Normalize operation on Medal is generated, the Normalize operation on Date is missing. This is because the phrase “the most” in the question suggests the need for type conversion of Medal, but there are no clues from the question to normalize Date, unless we observe from the table that the values are inconsistent.
@@
Key Challenge. The above example clearly demonstrates that the key challenge in designing the Planner agent, as discussed earlier, is to capture the relationships between different parts of question $Q$ and the columns in table $T$. For instance, as shown in Figure 4, the Normalize operation on the Date column is generated by considering two key factors: (1) the question contains “in February”, and (2) the values in the Date column are inconsistent.
@@ 5.2. Search Space of Function Pools
(3) $\mathtt{clean\mbox{-}string}$. To clean string values in tables, we define the function $\mathtt{clean\mbox{-}string}$. It has two arguments: ${\tt column}$ and ${\tt trans\_dict}$. The $\tt trans\_dict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ should be replaced with $v$.
(4) $\mathtt{infer}$. Similar to the one for the $\mathtt{Derive}$ operation, this function uses LLMs to deduce values that could not be processed by the functions above.
@@ 6.2. Improvement of Data Prep for TQA
Moreover, End2End and CoT methods also show considerable performance gains. These improvements are largely due to AutoPrep ’s filtering mechanism, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers from tables. Since “NL2SQL + AutoPrep” achieves the best accuracy in most cases, we take its results as default for further comparison.
@@ 6.5. In-Depth Exploration of System Capabilities
Exp-6: Evaluating extensibility on datasets requiring new logical operation types. To further evaluate the extensibility of AutoPrep, we evaluate it on a more complex dataset involving logical operations not originally supported. Since no existing TQA dataset presents sufficiently complex data quality issues, we construct a new dataset—TransTQ with the assistance of LLMs, which is available on our GitHub repository. Specifically, we prompt an LLM to identify tables from WikiTQ where inverse transformation operations (e.g., pivot and stack (Li et al., 2023)) can be applied. These operations are then executed to generate non-relational table formats, introducing new challenges for TQA on table transformation.
@@ 7. Related Work
However, previous TQA methods have the following limitations when handling data prep. First, all these methods do not address the column normalization tasks, which account for the most significant error types, as indicated in Figure 1. Second, all these methods utilize a single LLM agent for both data prep and answer reasoning. Given that data prep is a complex challenge, these one-size-fits-all solutions still fail to achieve satisfactory performance as demonstrated in our experiments.
Traditional Data Preparation Data prep techniques are widely used across various tasks, including training machine learning models for data analytics. Auto-Weka (Thornton et al., 2013) leverages Bayesian optimization to identify data prep operations. Auto-Sklearn (Feurer et al., 2020; Hutter et al., 2019) and TensorOBOE (Yang et al., 2020) apply meta-learning to discover promising operations. Alpine Meadow (Shang et al., 2019) introduces an exploration-exploitation strategy, while TPOT (Olson and Moore, 2016) uses a tree-based representation of data prep and genetic programming optimization techniques. Several studies (Berti-Équille, 2019; El et al., 2020; Heffetz et al., 2020) explore reinforcement learning techniques. HAIPipe (Chen et al., 2023) integrates both human-orchestrated and automatically generated data prep operations.
In this paper, we propose AutoPrep to generate question-aware data prep operations for TQA tasks. The key difference between AutoPrep and traditional data prep methods is that AutoPrep requires semantic alignment between the table and the NL question. Traditional data prep methods typically operate at an offline stage, independent of any specific downstream question or query. In contrast, question-aware data prep studied in AutoPrep tailors the tables to the specific needs of the question during the online stage, directly addressing the challenge of aligning the table’s structure with the NL question’s semantics.
@@
We adopt the hierarchical structure because our design introduces a Planner agent that decomposes the overall data preparation task into three distinct sub-tasks, each handled by specialized agents. This top-down coordination naturally aligns with the principles of the hierarchical multi-agent framework. Moreover, while AutoTQA focus on improving table analysis, this paper focuses on the performance bottleneck caused by data prep issues in TQA. Our proposed framework, AutoPrep, addresses question-aware data prep for TQA and can be integrated as a plugin into current TQA approaches to further improve the overall performance.
@@ 8. Conclusion and Future work
In this paper, we have introduced AutoPrep, an LLM-based multi-agent framework designed to support automatic data prep for TQA tasks. AutoPrep consists of three stages: (1) the Planning stage, which suggests logical data prep operations; (2) the Programming stage, which generates physical implementations for each logical operation; and (3) the Executing stage, which executes the Python code and reports error messages. We propose a Chain-of-Clauses method to generate high-quality logical plans and a Tool-augmented method for effective physical plan generation. Moreover, we introduce an effective method to ensure successful execution of the physical plan. Extensive experiments on WikiTQ, TabFact and TabBench demonstrate the superiority of AutoPrep.
@@
For future work, we identify three promising directions. First, we aim to extend the system’s data prep capabilities by integrating a broader set of operations to handle more complex issues such as missing values, duplicates, etc. Second, AutoPrep currently focuses on single-table question answering; extending its capabilities to support multi-table TQA is essential for tackling more realistic and complex scenarios. The third direction is to enable question-aware data preparation over enterprise datasets, which are much more complex than existing TQA benchmarks.
Appendix A Details of the Error Analysis
This section presents details of the error analysis shown in Figure 1, which motivates us to study question-aware data prep.
TQA Method: NL2SQL Based on GPT-4. We adopt NL2SQL (Rajkumar et al., 2022), a representative TQA baseline that does not perform data prep, as introduced in Section 6.1. Specifically, given a natural language question $Q$ and a table $T$, we prompt GPT-4 to generate an SQL query ${\tt SQL}$ using a few-shot in-context learning strategy. The query is then executed on $T$ to obtain the final answer $Ans$ for $Q$.
Error Categorization. We recruite three graduate students from our university’s database group to annotate error categories for a randomly sampled set of 500 instances from WikiTQ and TabFact. All annotators have a good understanding of the TQA task and are proficient in SQL syntax. We instruct the annotators to first identify errors stemming from insufficient data prep, assigning any remaining errors to a default category labeled Other. For the data prep-related errors, after careful investigation, we further categorize them into the following three types: (1) Missing Semantics: The table lacks necessary semantic information required to accurately answer the NL question. (2) Inconsistent Values: Values within a column are represented inconsistently across different records. (3) Irrelevant Columns: The table includes many columns, but only a small subset are actually relevant to the NL question.
In-depth Analysis of the “Other” Category. We report the detailed sub-categories within the “Other” category in Figure 9. As shown, for WikiTQ, more than half of the cases stem from issues inherent to the dataset itself. Notably, 40% of the errors are due to limitations in the evaluation function, where semantically equivalent but non-identical responses are marked as incorrect, a problem also highlighted in (Cheng et al., 2023). Additionally, 20% of the errors arise from formatting issues in the model’s responses, where the generated SQL logic is correct but the output includes extraneous values from irrelevant columns. On the other hand, for the TabFact dataset, most errors are attributed to model reasoning failures. Over 50% result from the model’s inability to generate logically complete and accurate SQL queries. Furthermore, more than 35% of the errors are due to response formatting issues, where the model returns ambiguous values instead of definitive true/false answers.
Appendix B More Experiments
B.1. Evaluation on Various Table Sizes
Exp: Evaluation of AutoPrep on tables with various sizes. To further investigate AutoPrep, we analyze its performance across tables of varying sizes. While all tables in TabFact are small, we categorize the tables in WikiTQ into Small (fewer than 2048 tokens), Medium (2048 to 4096 tokens), and Large (more than 4096 tokens), resulting in 4040 Small tables, 200 Medium tables, and 104 Large tables. The results are reported in Figure 10.
We observe that all baselines exhibit unstable performance, particularly on larger tables. For instance, although CoTable achieves the highest average accuracy among previous SOTA methods, it suffers an accuracy drop of 11.78 when reasoning over large tables. Likewise, ReAcTable shows relatively stable performance on medium tables (dropping by 2.15), yet its performance on large tables remains unsatisfactory (dropping by 17.53).
Figure 10. Accuracy on different size of tables in WikiTQ.
In contrast, AutoPrep achieves the highest and most stable performance across tables of varying sizes. When processing medium and large tables, the accuracy of AutoPrep drops by 2.29 and 3.79. The key to its stability is that it employs specialized agents for each data prep task, mitigating the issue of exceeding prompt length limits. Furthermore, each agent generates program-based operators to handle data prep on the entire table, thereby avoiding information loss that occurs when large tables are cropped.
Finding: By employing multiple agents and program-based operations, AutoPrep maintains stable performance as table size grows, ensuring that each data prep task is handled effectively without overwhelming a single model.
B.2. Evaluation on Various Question Difficulties
Table 6. In-depth Analysis of AutoPrep and all baselines when varying question difficulties on WikiTQ and TabFact.
Method WikiTQ TabFact
Simple Complex Simple Complex
End2End 57.8057.80{57.80}57.80 55.6855.68{55.68}55.68 $86.51$ $77.07$
CoT $56.85$ $55.33$ 84.9284.92{84.92}84.92 79.1379.13{79.13}79.13
NL2Py 61.8761.87{61.87}61.87 57.2157.21{57.21}57.21 $70.83$ $65.45$
NL2SQL $57.02$ $49.13$ $70.83$ $69.59$
ICL-Prep 59.5159.51{59.51}59.51 54.0254.02{54.02}54.02 $82.64$ 78.4478.44{78.44}78.44
Dater $51.33$ $45.77$ $88.39$ $77.76$
Binder $62.02$ $52.04$ $89.29$ $76.38$
ReAcTable 66.9866.98{66.98}66.98 61.7161.71{61.71}61.71 90.4890.48{90.48}90.48 79.7279.72{79.72}79.72
CoTable 66.8866.88{66.88}66.88 62.5362.53{62.53}62.53 90.5890.58{90.58}90.58 81.8981.89{81.89}81.89
AP + End2End $66.58$ ($+8.78$) $60.23$ ($+4.55$) $\underline{87.20}$ ($+0.69$) $77.07$ ($+0.00$)
+ CoT $64.12$ ($+7.27$) $58.57$ ($+5.24$) $84.52$ ($-0.40$) $80.02$ ($+0.89$)
+ NL2Py $\underline{68.09}$ ($+6.22$) 63.97 ($+6.76$) 90.77 ($+19.94$) $\underline{83.96}$ ($+18.51$)
+ NL2SQL 69.17 ($+11.97$) $\underline{63.50}$ ($+14.37$) 90.77 ($+19.94$) 84.94 ($+15.35$)
Also, following previous work (Ye et al., 2023; Wang et al., 2024), we compare AutoPrep and all baseline models on questions with different difficulties.
Exp: How does AutoPrep and other baselines perform on question with different difficulties? We divide the TabFact dataset based on officially provided question difficulty labels. Since WikiTQ does not provide difficulty labels, we divide the dataset based on question length following previous work (Ye et al., 2023).
As shown in Table 6, CoTable achieves the best performance among all baselines on both Simple and Complex questions, indicating its abilities of complex reasoning abilities over tables. While AutoPrep still outperforms CoTable, especially on Complex questions. Specifically, AutoPrep outperforms CoTable by 0.21% on Simple questions and 3.72% on Complex questions.
Moreover, notice that trivial performance improvement is achieved by AutoPrep for Direct Prompting methods. Especially for CoT on Simple questions, the performance of CoT has been dropped by 0.47% when integrated with AutoPrep. This is because tables of TabFact are all small tables, data prep operations of AutoPrep fail to bring obvious benefit for Direct Prompting methods. While for Code Generation methods, data quality is a critical problem. Thus, NL2Code and NL2SQL have achieved obvious performance gains on prepared tables.
B.3. Impact of LLM-Based $\mathtt{CustomCode}$
Table 7. Analyzing LLM generated code for logical operator.
Method  WikiTQ
$\mathtt{CustomCode}$ ratio ${1.08\%}$
$\mathtt{CustomCode}$ accuracy ${55.32\%}$
AutoPrep w. $\mathtt{CustomCode}$ 66.09
AutoPrep w.o. $\mathtt{CustomCode}$ 65.6365.63{65.63}65.63
Exp : How does LLM-based $\mathtt{CustomCode}$ help AutoPrep generalize to cases requiring new physical operations? As discussed in Section 5.1, we prompt the LLM to write code to implement the logical operation if no suitable function in the function pool can meet the requirement, which we denote as $\mathtt{CustomCode}$. We quantify the usage frequency of $\mathtt{CustomCode}$ and evaluate whether $\mathtt{CustomCode}$ produces the correct code on WikiTQ. We also evaluate the performance of AutoPrep without $\mathtt{CustomCode}$ design.
As shown in Table 7, The results show that $\mathtt{CustomCode}$ is invoked in 1.08% of cases, achieving an accuracy of 55.32%. Notably, removing $\mathtt{CustomCode}$ leads to a performance drop of 0.46 points (from 66.09 to 65.63). These results indicate that, despite its relatively low invocation frequency and moderate accuracy, $\mathtt{CustomCode}$ improves generalization of AutoPrep, particularly in handling new or unforeseen physical operations.
Finding : $\mathtt{CustomCode}$ improves the generalization of AutoPrep on cases requiring new physical operations.
B.4. Analysis on Other Data Quality Issues
We have conducted a new experiment to analyze the impact of missing values and duplication in TQA tasks on the WikiTQ dataset.
Specifically, for duplication, we first examine all tables containing duplicated records using a heuristic method from (Fan et al., 2024) on the WikiTQ dataset, and only obtain 24 such instances. Then, we manually remove the duplicate records and re-evaluate AutoPrep ’s performance on the updated dataset. The results remain unchanged (accuracy: 70.83% for both the original and de-duplicated datasets), suggesting that duplication is not a critical issue on existing TQA benchmarking datasets. Similarly, for missing values, we prompt LLMs to impute the missing entries based on the available data for all tables and then re-evaluate AutoPrep on the imputated dataset. The results also show insignificant difference (65.00% w/o imputation vs. 64.33% with imputation), indicating that missing values have limited impact.
These above results suggest that such data preparation issues are uncommon and have little effect on existing TQA benchmarking datasets. This is largely due to the way of constructing the datasets. Specifically, during dataset construction (Chen et al., 2019; Pasupat and Liang, 2015; Wu et al., 2024), annotators are instructed to first generate natural language questions based on a given table and then derive the corresponding answers directly from the same table. This construction paradigm implicitly ensures data completeness and consistency, as annotators tend to avoid crafting questions around missing or duplicating information. As a result, common data quality issues such as missing values and duplicate records are naturally excluded from the dataset, making them less relevant in the context of existing TQA tasks.
Nevertheless, we acknowledge that issues such as missing values and duplicates are common in real-world datasets, we leave the exploration of these broader data prep challenges to future work.
Finding : While issues such as missing values and duplicates are common in real-world datasets, they are significantly less prevalent than the three highlighted challenges in existing TQA benchmarks.
Appendix C A Running Example of AUTOPREP
(a) Case one
(b) Case two
Figure 11. Two case studies selected from TabBench and WikiTQ to qualitatively analyze effectiveness of AutoPrep.
This section presents two illustrative examples from TabBench and WikiTQ to qualitatively analyze effectiveness of AutoPrep.
To answer the question in Figure 11(a), CoTable generates an operator chain with an “add_col” operator, which is used to generate a new column Tol based on the division results of two existing columns Expense and Percent. This operator directly prompts an LLM to output a list containing all values of the new column, which is a challenging task for previous LLMs. Thus, CoTable generates a new column with two wrong values which leads to a wrong answer. For ReAcTable, although it generates a logically correct Python program, it ignores the types of existing columns which do not support for numerical calculation. Thus, ReAcTable also fails to derive the Tol column. AutoPrep addresses this issue by first utilizing a Planner agent to generate logical operators specifying data prep requirements, i.e., two $\mathtt{Normalize}$ to normalize column Expense and Percent, a $\mathtt{Derive}$ to generate the overall expenses Tol and a $\mathtt{Filter}$ to select related columns Year and Tol. These logical operators are passed to Programmer agents to generate physical operators consisting of a pre-defined Python function. For example, we call $\mathtt{to\mbox{-}numerical}$ function two times to cast Expense and Percent to numerical values. All physical operators are executed to process the original table and output a prepared table. Finally, an Analyzer agent based on NL2SQL method generates a SQL query to extract the correct answer “4214.17”.
Figure 11(b) illustrates a TQA instance involving large tables and inconsistent values. AutoPrep solves this by splitting the TQA question into two phases, i.e., data prep and data analysis. It first generates logical operators including two $\mathtt{Normalize}$ and one $\mathtt{Filter}$. Next, the Programmer agents implement them with corresponding physical operators, which are executed to produce a prepared table. Based on this, the Analyzer agent extracts the final answer “8”.
Appendix D Prompts in AutoPrep
The representative demonstration examples and prompts used by AutoPrep are presented.
D.1. Prompts in Planner Agent
Figure 12. Prompt of Planner on generating Analysis Sketch on WikiTQ Dataset.
Figure 13. Prompt of Planner on generating Analysis Sketch on TabFact Dataset.
Figure 14. Prompt of Planner on generating requirement for Derive Operator.
Figure 15. Prompt of Planner on generating requirement for Normalize Operator.
The Planner aims at generating logical operators in two stages: (1) generate the QA-Sketch. (2) generate specific requirement for Normalize and Derive.
In the first stage, AutoPrep analyzes the table and the question to generate the reasoning process of question answering over table with the assistance of pseudo codes. In this paper, we define the grammar of the pseudo code as SQL with a UDF grammar. The prompts for WikITQ and TabFact datasets are shown in Figure 12 and Figure 13. Next, based on the generated QA-Sketch, we extract each operation clause and provide the related column data in the clause to query LLMs to generate specific requirement of the logical operator. The prompts of generating specific requirement for each type of logical operator are shown in Figure 14 and Figure 15. Notice that the Filter operator is generated by analyzing the QA-Sketch and the arguments of Derive operator, which does not need to query LLMs.
D.2. Prompts in Programmer Agents
Figure 16. Prompt of Derive.
Figure 17. Prompt of Normalize.
Figure 18. Prompt of Imputater.
Figure 19. Prompt of LLM Coder.
Given logical operators with specific requirement and related columns, the programmer agents generate physical operators to implement the logical operators. The prompts of Derive and Normalize are shown in Figure 16 and Figure 17. In addition, if any logical operators can not be implemented with functions in the function pool, we call LLMs to generate Python codes. The prompt of generating Python code is shown in Figure 19. Finally, if any values can not be derived or normalized, we call an Imputater agent to direct infer the correct values. The prompt is shown in Figure 18.
D.3. Prompts in other baselines
Figure 20. Prompt of End2End method on WikiTQ dataset.
Figure 21. Prompt of End2End method on TabFact dataset.
Figure 22. Prompt of Chain-of-Thought TQA method on WikiTQ dataset.
Figure 23. Prompt of Chain-of-Thought TQA method on TabFact dataset.
Figure 24. Prompt of NL2SQL on WikiTQ dataset.
Figure 25. Prompt of NL2SQL on TabFact dataset.
Figure 26. Prompt of ICL-Prep on WikiTQ dataset.
Figure 27. Prompt of ICL-Prep on TabFact dataset.
Figure 28. Prompt of NL2Py on WikiTQ dataset.
Figure 29. Prompt of NL2Py on TabFact dataset.
We also provide the prompts of implemented baselines, including End2End (Figure 20 and Figure 21), CoT (Figure 22 and Figure 23), NL2SQL (Figure 24 and Figure 25), NL2Py (Figure 28 and Figure 29) and ICL-Prep (Figure 26 and Figure 27). Some of the prompts are from existing works (Cheng et al., 2023; Chen, 2022) as described in the experimental setup.

### v4

@@ Abstract.
Proceedings of the VLDB Endowment, Vol. 18, No. 10 ISSN 2150-8097.
@@
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/ruc-datalab/AutoPrep.
@@ 1. Introduction
Tabular Question Answering (TQA) refers to the task of answering natural language (NL) questions based on provided tables (Lu et al., 2025; Cheng et al., 2022, 2023; Pal et al., 2023). TQA empowers non-technical users such as domain scientists to easily analyze tabular data and has a wide range of applications, including table-based fact verification (Chen et al., 2019; Gu et al., 2022a, b) and table-based question answering (Pasupat and Liang, 2015; Nan et al., 2022). As TQA requires NL understanding and reasoning over tables, state-of-the-art solutions (Wei et al., 2022; Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024; Wang et al., 2024) mainly rely on large language models (LLMs).
@@
(3) Irrelevant Columns. In some tables, although many columns are available, only a few are actually relevant to the NL question. For example, as shown in Figure 2(c), $Q_{7}$ does not require the Summary column in table $T_{7}$. Similarly, in $T_{8}$, although there are 12 columns, only 2 are needed to answer $Q_{8}$. Thus, column filtering, i.e., retaining only the columns directly related to the question, is crucial for TQA, as irrelevant columns add complexity and can mislead LLMs into producing incorrect answers.
This highlights the critical role of thorough data prep in ensuring accurate TQA. Note that, although other issues like missing values and duplicates are common in real-world data, they are less prevalent than the three highlighted issues in current TQA benchmarks. Nevertheless, we recognize their significance and leave the exploration of broader data prep challenges to future work.
@@
AutoPrep: A Hierarchical, Multi-Stage Approach. To address the challenges, we propose AutoPrep, which features two key ideas. First, drawing inspiration from modern DBMS, particularly the distinction between logical operations and their physical implementations, AutoPrep separates high-level, logical data prep operations from the concrete methods used for execution. Specifically, it introduces a planning stage that generates a logical plan for each question, consisting of a sequence of high-level data prep operations tailored to the question’s needs, such as column derivation, normalization and filtering shown in Figure 2. In the next stage, AutoPrep maps these logical operations to the corresponding physical implementations. This separation allows AutoPrep to break down complex data prep tasks into smaller, more manageable sub-tasks, making the process easier to solve. Moreover, this modular design makes AutoPrep extensible, enabling the development of specialized implementations for each type of data prep operation or the introduction of new operation types.
@@ 2.1. Tabular Question Answering
Figure 2 provides several examples of TQA. Consider table $T_{1}$, which contains medal information for cyclists from different countries, with two columns: Cyclist and Medal. Given the question $Q_{1}$, “Which country has the most medals?”, the answer should be ITA, as two Italian cyclists, “Dav” and “Alex”, have won a total of 4 medals, more than the ESP cyclist “Alej”. TQA often requires reasoning over tables. For instance, to answer question $Q_{2}$, we first need to calculate the “GDP growth rate” for all countries, then sort the countries by growth rate, and finally identify the country with the highest GDP growth rate, i.e., CHN.
@@ 2.2. Data Prep for TQA
(a) The Derive operation: Figure 2(a) shows three examples of the Derive operations, i.e., extracting Country information from column Cyclist in $T_{1}$, computing new column GrowthRate using two columns in $T_{2}$, and inferring a status of IfLost by analyzing the scores in column Result.
@@ 2.3. A Straightforward LLM-based Solution
First, at the logical-operation level, given the inherent difficulties in understanding both NL questions and tables, it is challenging to accurately identify which data prep operations are specifically required to satisfy the needs of the NL question. This often leads to false negatives and false positives. For example, as shown in Figure 3, converting the Age column to a numerical format in code block ② is a false positive, as it is irrelevant to the question. In contrast, failing to normalize the Date column before extracting the month in code block ③ constitutes a false negative, as the method ignores the inconsistency in Date formats.
@@ 3. An Overview of AutoPrep
    Multiple Programmer Agents (e.g., Normalize): the Programming stage. It directs the LLM to generate physical implementation $f_{i}$ (e.g., Python code) for each operation $o_{i}$ customized for the table $T$. Besides, it is also tasked for code debugging if any execution errors occur.
@@
  • Normalize(“Format date as %m-%d”, Date) standardizes the values in the Date column into a unified format to support the “in February” condition in the question.
@@
(b) The Programming stage: AutoPrep designs specialized Programmer agents for each operation type, i.e., Derive, Normalize and Filter. Each specialized Programmer focuses on generating executable code for its assigned operations.
@@
The Planner Agent. The key challenge is how to suggest a logical plan that address specific NL questions. Even for the same table, different NL questions may require not only different logical operations but also varying sequences of those operations. To address this challenge, we propose a novel Chain-of-Clauses (CoC) reasoning method for the Planner agent. This method translates the NL question into an Analysis Sketch, outlining how the table should be transformed to produce the answer, thereby guiding the agent’s reasoning based on this sketch. More details of the method are given in Section 4.
@@ 4.1. A Direct Prompting Method
(b) Missing operations: Observing the ground-truth in Figure 4, we can see that the Derive operation on Cyclist is not generated, as the column has already been filtered out. Moreover, although the Normalize operation on Medal is generated, the Normalize operation on Date is missing. This is because the phrase “the most” in the question suggests the need for type conversion of Medal, but there are no clues from the question to normalize Date, unless we observe from the table that the values are inconsistent.
@@
The above example clearly demonstrates that the key challenge in designing the Planner agent, as discussed earlier, is to capture the relationships between different parts of question $Q$ and the columns in table $T$. For instance, as shown in Figure 4, the Normalize operation on the Date column is generated by considering two key factors: (1) the question contains “in February”, and (2) the values in the Date column are inconsistent.
@@ 5.2. Search Space of Function Pools
(3) $\mathtt{clean\mbox{-}string}$. To clean string values in tables, we define function $\mathtt{clean\mbox{-}string}$ with two arguments: ${\tt column}$ and ${\tt trans\_dict}$. The $\tt trans\_dict$ is a dictionary where each key $k$ and its corresponding value $v$ represent that $k$ in $\tt column$ is replaced by $v$.
(4) $\mathtt{infer}$. This function uses LLMs to deduce values that could not be processed by the functions above.
@@ 6.2. Improvement of Data Prep for TQA
Moreover, End2End and CoT methods also show considerable performance gains. These improvements are largely due to the filtering mechanism of AutoPrep, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers from tables. Since “NL2SQL + AutoPrep” achieves the best accuracy in most cases, we take its results as default for further comparison.
@@ 6.5. In-Depth Exploration of System Capabilities
Exp-6: Evaluating extensibility on datasets requiring new logical operation types. To further evaluate the extensibility of AutoPrep, we evaluate it on a more complex dataset involving logical operations not originally supported. Since no existing TQA dataset presents sufficiently complex data quality issues, we construct a new dataset TransTQ with the assistance of LLMs, which is available on our GitHub repository. Specifically, we prompt an LLM to identify tables from WikiTQ where inverse transformation operations (e.g., pivot and stack (Li et al., 2023)) can be applied. These operations are then executed to generate non-relational table formats, introducing new challenges for TQA on table transformation.
@@ 7. Related Work
However, previous TQA methods do not address the column normalization tasks, which account for the most significant error types, as indicated in Figure 1. Second, the methods utilize a single LLM agent for both data prep and answer reasoning. Given that data prep is a complex challenge, these one-size-fits-all solutions may not achieve satisfactory performance.
Traditional Data Preparation. Data prep techniques are widely used across various tasks (Chai et al., 2023; Fan et al., 2024; Tu et al., 2022). For training machine learning models for data analytics, Auto-Weka (Thornton et al., 2013) leverages Bayesian optimization to identify data prep operations. Auto-Sklearn (Feurer et al., 2020; Hutter et al., 2019) and TensorOBOE (Yang et al., 2020) apply meta-learning to discover promising operations. Alpine Meadow (Shang et al., 2019) introduces an exploration-exploitation strategy, while TPOT (Olson and Moore, 2016) uses a tree-based representation of data prep and genetic programming optimization techniques. Several studies (Berti-Équille, 2019; El et al., 2020; Heffetz et al., 2020; Zhang et al., 2025) explore reinforcement learning techniques. HAIPipe (Chen et al., 2023) integrates both human-orchestrated and automatically generated data prep operations.
We propose AutoPrep to generate question-aware data prep operations for TQA tasks. Traditional data prep methods typically operate at an offline stage, independent of any specific downstream question or query. In contrast, question-aware data prep studied in AutoPreptailors the tables to the specific needs of the question during the online stage, directly addressing the challenge of aligning the table’s structure with the NL question’s semantics.
@@
We adopt the hierarchical structure because our design introduces a Planner agent that decomposes the overall data preparation task into three distinct sub-tasks, each handled by specialized agents. This top-down coordination naturally aligns with the principles of the hierarchical multi-agent framework. Moreover, while AutoTQA focuses on improving table analysis, this paper focuses on the performance bottleneck caused by data prep issues in TQA. Our proposed framework, AutoPrep, addresses question-aware data prep for TQA and can be integrated as a plugin into current TQA approaches to further improve the overall performance.
@@ 8. Conclusion and Future work
In this paper, we have introduced AutoPrep, an LLM-based multi-agent framework to support data prep for TQA tasks. AutoPrep consists of three stages: (1) the Planning stage, which suggests logical data prep operations, (2) the Programming stage, which generates physical implementations for each logical operation, and (3) the Executing stage, which executes the Python code and reports error messages. We propose a Chain-of-Clauses method to generate high-quality logical plans and a Tool-augmented method for effective physical plan generation. Extensive experiments on real datasets demonstrate the superiority of AutoPrep.
@@
For future work, we identify three promising directions. First, we aim to extend the system’s data prep capabilities by integrating a broader set of operations to handle more complex issues such as missing values, duplicates, etc. Second, AutoPrep currently focuses on single-table question answering, and extending its capabilities to support multi-table TQA is essential for tackling more realistic and complex scenarios. The third direction is to enable question-aware data preparation over enterprise datasets, which are more complex than existing TQA benchmarks.
Acknowledgment
This paper was partly supported by the NSF of China (62436010, 62441230, 62525202 and 62232009), National Key R&D Program of China (2023YFB4503600), NSF DBI-2327954, Guangdong Provincial Project 2023CX10X008 and Amazon Research Awards.